# Binaural Target Speaker Extraction using HRTFs and a Complex-Valued Neural Network

### v1

@@ Abstract
In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers. To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker. Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets in different languages. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals. This deviates from conventional approaches that use spectrograms or treat the real and imaginary components of the STFT as separate real-valued inputs. We begin by evaluating the method in an anechoic, noise-free scenario, where it demonstrates excellent extraction performance while effectively preserving the binaural cues of the target signal. Next, we test a modified variant under mild reverberation conditions. This version remains robust in reverberant conditions, maintaining speech clarity, preserving source directionality, and simultaneously reducing reverberation. Demo-page: https://bi-ctse-hrtf.github.io
@@ I Introduction
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims at selectively extracting one desired speaker of the multiple competing speakers, in a binaural setting. In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids feature multiple microphones on each earpiece and fall into one of the two categories: 1) Bilateral systems, in which each device operates independently, yielding a single-channel output. 2) Shared systems, where both earpieces share all microphones, producing a dual-channel output.
@@
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, Target Speaker Extraction (TSE) algorithms leverage prior knowledge of the target speaker(s) to isolate their voice. The prior knowledge about the speaker—commonly referred to as “clue” [6]—can include any additional information about the target speaker. For instance, it might be visual data depicting the speaker’s appearance as in [7] that uses a face image to extract its corrospondig speech, or [8] that uses both audio and visual clues, where the audio clue is an enrollment utterance and the visual clue is a video of the speaker speaking within the mixture. Another popular clue is a pre-recorded sample used to extract speaker-specific features, as in [9, 10, 11, 12].
While examining prior works, such as [13], we observed the particular challenge of separating two voices with highly similar characteristics—for example, speakers of the same gender. Consequently, we opted for an approach that does not depend on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA. These spatially guided methods consistently outperform single-microphone solutions.
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that leverages the listener’s HRTF, aligned with the target speaker’s DOA, as an extraction clue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. Unlike real‑valued models that treat real and imaginary parts separately—or architectures that rely solely on the magnitude spectrogram—our network processes the full complex representation, more effectively preserving phase information. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases.
@@ II Problem Formulation
Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation). Similarly, the interfering speaker’s DOA is ($\theta_{i}$,$\phi_{i}$). Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively. This signal can be written as
@@
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech of the desired speaker and the interfering speaker, respectively. Let $\gamma=\{d,i\}$ denote the speaker index, either the desired ($d$) or the interferer ($i$). Here, $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ is a Transfer Function (TrF) modeling acoustic propagation from source $\gamma$ to both earpieces as a superposition of HRTFs associated with the DOAs of all reflections. The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, represents the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$. According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$ Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
Our work has three main objectives: 1) to extract the target speaker at a known DOA from binaural input using the listener’s HRTF; 2) to generate a binaural output that preserves the spatial cues of the desired source; and 3) to dereverberate the output, retaining only the direct-path HRTF.
@@ III Proposed Method
We propose a deep neural network model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task. Given the HRTF from the desired source DOA $(\theta_{d},\phi_{d})$, namely $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. We denote the extracted signal as $\hat{\tilde{\bm{s}}}_{d}$.
@@ III-A Architecture
The proposed model architecture is based on the approach presented in [14], modified to fully handle complex-valued inputs, rather than the real and imaginary components independently.
@@
To capture cross-channel (left and right) information in the HRTF, we apply the self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage. A block diagram of this architecture is presented in Fig. 1.
@@
Figure 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed signal at the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of estimated desired signal.
@@ III-B Loss function
[…]
@@
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a two-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality.
@@ IV-A Data Generation
The RIEC HRTF Dataset [19, 20]: A measured HRTF database consists of 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
@@
Figure 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$ [^2] Images sourced from: https://www.freepik.com
We employed the SofaMyRoom framework [23] to simulate reverberant and anechoic conditions, similar to (6). This tool places a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a target reverberation time. It also supports zero-order simulations for anechoic scenarios. For training, we sampled $T_{60}\sim\mathcal{U}[0.2,0.8]\,$s, and for testing, $T_{60}\sim\mathcal{U}[0.1,0.4]\,$s. Speaker and listener positions were also drawn uniformly within the room’s dimensions. The simulation arrangement is illustrated in Fig. 2.
IV-B Algorithm Setup
The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, and transformed using a 512-sample STFT with 75% overlap. Our model was trained in a supervised manner. For the anechoic setting, the target signals were generated by convolving the desired speech signal with the corresponding HRTF. For the reverberant setting, the targets were obtained by simulating a reverberant room and using only the first arrival (i.e., zero-order reflection). The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and transformed to the frequency domain using Fast Fourier Transform (FFT).
@@ IV-C Metrics
To evaluate our method’s performance, we used SI-SDR[27], Perceptual Evaluation of Speech Quality (PESQ)[28], and Short-time Objective Intelligibility (STOI)[29] as measures of speech quality and intelligibility. lFurthermore, to verify that the estimated signals retain their binaural characteristics, we compare the Probability Density Function (p.d.f.) of the ILD and ITD of the target signals and the extracted signals. The similarity of the p.d.f.s indicates that the spatial cues are preserved.
@@ IV-D Competing Methods
To ensure a fair comparison, all baseline methods operate under an identical setup, namely, binaural TSE using only two microphones (one per ear). Accordingly, we compare our method against the BLCMV beamformer [4] and a recent DNN-based model [30], both designed specifically for this setup. In addition, we trained two variants of our model. The first is Bi-TSE-HRTF-RI, a real-valued model, in which the complex-value input features are split into Real-Imaginary (RI) components and concatenated, resulting in four input channels instead of the two input channels in the original scheme. The same procedure is applied to the HRTF clue. The second is Bi-TSE-DOA, a complex-valued model that incorporates the DOA as a clue instead of the HRTF. Here, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs and passed through a fully-connected layer to match the dimensionality of the bottleneck. For the specified variants, both terms of (5) remain intact, i.e., the target signal is the binaural rendering of the desired speech signal with the correct HRTF.
(a) SI-SDR[dB]
(b) PESQ
(c) STOI
Figure 3: Histograms of speech quality and intelligibility measures, before and after the application of the proposed TSE algorithm for the anechoic setting.
TABLE I: The evaluation datasets include 3k speech samples mixed with SIR sampled from $\mathcal{U}[0,5]$ dB. Speaker overlap is either sampled from $\mathcal{U}[0\%,100\%]$ (designated by) or set to 100% (designated by∗∗), with audio lengths of 5 s and 8 s, respectively. Best scores per metric and dataset are highlighted in bold.
Method Dataset Reverberant SI-SDR (dB) $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ $\Delta$ILD [dB] $\downarrow$ $\Delta$ITD [mS] $\downarrow$
Unprocessed LS∗∗ 2.4 1.36 0.78 - -
BLCMV Beamformer [4] LS∗∗ 0.86 3.93 0.95 0.7341 0.6032
Bi-CSim-TSE [30] LS 18.52 3.05 0.96 - -
Bi-TSE-HRTF (proposed) LS∗∗ 19.35 3.27 0.98 0.1311 0.0033
Unprocessed MLS-French∗∗ 2.37 1.29 0.77 - -
BLCMV Beamformer [4] MLS-French∗∗ 0.98 3.92 0.95 0.7092 0.5829
Bi-TSE-HRTF (proposed) MLS-French∗∗ 18.9 3.23 0.98 0.2697 0.0120
Unprocessed WSJ0∗∗ 0.06 1.23 0.73 - -
BLCMV Beamformer [4] WSJ0∗∗ 1.51 3.98 0.96 0.733 0.5669
Bi-TSE-DOA WSJ0∗∗ 16.3 2.47 0.97 0.1304 1.0145
Bi-TSE-HRTF-RI WSJ0∗∗ 21.36 2.97 0.99 0.0934 0.00435
Bi-TSE-HRTF (proposed) WSJ0∗∗ 20.37 3.49 0.99 0.1003 0.0027
Unprocessed WSJ0∗∗ -1.15 1.21 0.7 - -
BLCMV Beamformer [4] WSJ0∗∗ 1.37 1.6 0.82 0.3303 0.5439
Bi-TSE-DOA WSJ0∗∗ 9.53 2.08 0.93 0.2410 1.0125
Bi-TSE-HRTF (proposed) WSJ0∗∗ 10.37 2.60 0.95 0.1998 0.0036

### v2

@@ Abstract
In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers. To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker. Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets and languages. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals, and compare it to a Real-Imaginary (RI)-based neural network, demonstrating the advantages of the former. We first evaluate the method in an anechoic, noise-free scenario, achieving excellent extraction performance while preserving the binaural cues of the target signal. We then extend the evaluation to reverberant conditions. Our method proves robust, maintaining speech clarity and source directionality while simultaneously reducing reverberation. A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods demonstrates that our approach attains performance on par with competing techniques in terms of noise reduction and perceptual quality, while offering a clear advantage in preserving binaural cues. Demo-page: https://bi-ctse-hrtf.github.io
@@ I Introduction
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims to selectively extract the desired speaker from multiple competing speakers in a binaural setting. In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids employ either a single microphone or multiple microphones per earpiece and can be broadly categorized into two groups: 1) Bilateral systems, where each device operates independently and provides a single-channel output per earpiece; 2) Shared systems, where all microphones across both earpieces are jointly utilized, yielding a dual-channel output.
@@
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge of the target speaker(s) to isolate their voice. This prior knowledge—commonly referred to as a “clue” [6]—can take various forms. A widely used clue is a pre-recorded speech sample, often referred to as an enrollment utterance, which provides speaker-specific features [7, 8, 9, 10]. Clues may also come from other modalities, such as visual data. For example, [11] uses a face image to extract the corresponding speech, while [12] combines audio and visual clues, with the audio clue being an enrollment utterance and the visual clue a video of the speaker speaking within the mixture.
While examining prior works, such as [13], we observed a particular challenge in separating two voices with highly similar characteristics, for example, speakers of the same gender. Consequently, we opted for an approach that does not depend on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA and extracts spatial features from it. These spatially guided methods consistently outperform single-microphone solutions.
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction cue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
@@ II Problem Formulation
Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation). Similarly, the interfering speaker’s DOA is ($\theta_{i}$$\phi_{i}$). Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively. The binaural signal can be written as
@@
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech signals of the desired speaker and the interfering speaker, respectively. Let $\gamma\in\{d,i\}$ denote the speaker index, where $d$ refers to the desired source and $i$ to the interferer. The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all reflections. The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizes the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$.
According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$. Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
Given the HRTF corresponding to the desired source’s DOA $(\theta_{d},\phi_{d})$, namely the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. The extracted signal is denoted by $\hat{\tilde{\bm{s}}}_{d}$. This binaural output is required to: 1) preserve the spatial cues of the desired source, and 2) dereverberate the output by retaining only the direct-path HRTF, thereby also dereverberating the desired speech signal.
@@ III Proposed Method
We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
@@ III-A Architecture
The proposed model architecture is based on the approach presented in [14], modified to fully handle dual-channel, complex-valued inputs, rather than the real and imaginary components independently. The proposed model employs a U-Net architecture, incorporating the HRTF at the bottleneck. The individual modules comprising the architecture are described in detail below.
@@
To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage. A block diagram of the proposed architecture is presented in Fig. 1.
@@
Fig. 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated binaural desired signal.
@@ III-B Loss function
$\text{SI-SDR}(\bm{s},\hat{\bm{s}})=20\log_{10}\left(\frac{\|\beta\bm{s}\|}{\|\beta\bm{s}-\hat{\bm{s}}\|}\right);\quad\beta=\frac{\hat{\bm{s}}^{\top}\bm{s}}{\|\bm{s}\|^{2}}\,$, (3)
@@
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality. The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
@@ IV-A Data Generation
The RIEC HRTF Dataset [19, 20]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
We employed the SofaMyRoom framework [23] to simulate both anechoic and reverberant conditions, following the formulation in (6). This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time. It also supports zero-order simulations for anechoic scenarios. The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room. The overall simulation setup is illustrated in Fig. 2.
IV-B Algorithm Setup
The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, then transformed with a 512-sample STFT (75% overlap). The model is trained in a supervised manner: in the anechoic setting, targets are generated by convolving the desired speech with the corresponding HRTF; in the reverberant setting, by simulating a room and keeping only the first arrival (zero-order reflection). The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and converted to the frequency domain via Fast Fourier Transform (FFT).
@@ IV-B Algorithm Setup
Fig. 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$. Images source: freepik.com
@@ IV-C Metrics
To evaluate our method, we employed SI-SDR [27], wide-band and narrow-band Perceptual Evaluation of Speech Quality (PESQ[28], and Short-time Objective Intelligibility (STOI[29] as intrusive metrics of speech quality and intelligibility, i.e., measures that compare the extracted signal against a reference target. In addition, we used DNS-MOS [30], a non-intrusive DNN-based predictor of Mean Opinion Score (MOS), which estimates speech quality (SIG), background noise quality (BAK), and overall quality (OVRL).
[…]
(a) Reverberant Mixture
(b) $\tilde{\bm{s}}_{d}$
(c) $\hat{\tilde{\bm{s}}}_{d}$
(d) $\tilde{\bm{s}}_{i}$
(e) $\hat{\tilde{\bm{s}}}_{i}$
Ground-truthBi-TSE-HRTF (proposed)BDE-BiTSE[15]BLCMV Beamformer [4]
Fig. 3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced. $\tilde{\bm{s}}_{d}$ and $\tilde{\bm{s}}_{i}$ denotes the anechoic target signals. Graphs produced by [31].
Table I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples. Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap . Best scores are highlighted in bold.
[…]
LS Unprocessed (2.688\3.380\3.140) -0.01 1.23 1.61 0.72 - -
BLCMV Beamformer [4] (2.827\3.304\3.495) -1.0 3.46 3.9 0.95 0.672 0.599
BDE-BiTSE [15] (3.072\3.471\3.768) 24.01 3.66 4.02 0.99 0.142 0.485
Bi-TSE-HRTF (proposed) (2.999\3.398\3.742) 20.18 3.14 3.78 0.98 0.1024 0.0042
MLS-French Unprocessed (2.499\3.252\2.947) -0.01 1.23 1.63 0.71 - -
BLCMV Beamformer (2.761\3.277\3.384) -1.03 3.37 3.82 0.94 0.694 0.608
BDE-BiTSE (2.987\3.403\3.709) 23.67 3.67 4.04 0.99 0.3 0.503
Bi-TSE-HRTF (proposed) (2.900\3.324\3.655) 19.86 3.14 3.79 0.98 0.1039 0.0041
WSJ0 Unprocessed (2.845\3.376\3.411) 0 1.25 1.57 0.72 - -
BLCMV Beamformer (3.044\3.444\3.748) -0.42 3.48 3.84 0.95 0.745 0.613
BDE-BiTSE (3.230\3.543\3.951) 25.59 3.77 4.01 1.0 0.126 0.484
Bi-TSE-DOA (2.965\3.376\3.660) 16.44 2.43 3.09 0.97 0.165 0.990
Bi-TSE-HRTF-RI (3.110\3.504\3.764) 21.39 2.93 3.57 0.99 0.085 0.0044
Bi-TSE-HRTF (proposed) (3.18\3.501\3.911) 21.41 3.47 3.88 0.99 0.0035 0.0756
WSJ0 (Reverb) Unprocessed (2.348\3.069\2.783) -1.9 1.15 1.44 0.69 - -
BLCMV Beamformer (1.840\2.612\2.038) -1.11 1.35 1.88 0.74 0.430 0.593
BDE-BiTSE (3.141\3.438\3.952) 14.8 3.09 3.56 0.98 0.204 0.386
Bi-TSE-DOA (2.681\3.196\3.324) 8.29 1.82 2.44 0.91 0.296 0.990
Bi-TSE-HRTF-RI (3.030\3.362\3.838) 11.18 2.01 2.74 0.95 0.174 1.001
Bi-TSE-HRTF (proposed) (2.86\3.348\3.481) 10.67 2.17 2.79 0.94 0.186 0.0082
@@ IV-D Competing Methods
To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear). Accordingly, we compared our method against the BLCMV beamformer [4] and a recent DNN-based model [15], both specifically designed for this scenario. The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the same training data as ours. We ensured that the strongest variant reported in the paper (BDE+CDF+IPD+SDF) was used for comparison.
In addition, we trained two variants of our model. The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are split into their RI components and concatenated, resulting in four input channels instead of two. The same decomposition is applied to the HRTF clue. The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue instead of the HRTF. In this case, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs, and passed through a fully connected layer to match the dimensionality of the bottleneck. For both variants, the loss function in (5) remains unchanged, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.